# A Frank-Wolfe-based primal heuristic for quadratic mixed-integer optimization

### v1

@@ Abstract
We propose a primal heuristic for quadratic mixed-integer problems. Our method extends the Boscia framework – originally a mixed-integer convex solver leveraging a Frank-Wolfe-based branch-and-bound approach – to address nonconvex quadratic objective and constraints. We reformulate nonlinear constraints, introduce preprocessing steps, and a suite of heuristics including rounding strategies, gradient-guided selection, and large neighborhood search techniques that exploit integer-feasible vertices generated during the Frank-Wolfe iterations. Computational results demonstrate the effectiveness of our method in solving challenging MIQCQPs, achieving improvements on QPLIB instances within minutes and winning first place in the Land-Doig MIP Computational Competition 2025.
@@ 1 Introduction
Mixed-Integer Quadratically Constrained Quadratic Problems (MIQCQPs) represent a broad category in optimization, capturing an array of applications in operations research and machine learning. These problems combine the combinatorial difficulties of mixed-integer structures with nonlinear nonconvex constraints. Despite this double difficulty, solution methods, algorithms, and solvers have been developed over the past decades for generic or specific forms of MIQCQPs, including Couenne [2], GloMIQO [28], ANTIGONE [29], BARON [34], and SCIP [8, 10, 9]. We refer the reader to [24, 25] for recent reviews of mixed-integer nonlinear optimization beyond quadratic functions. Recently, commercial mixed-integer linear solvers such as Xpress and Gurobi added capabilities to solve MIQCQPs to global optimality, showing the strong interest from application areas. We denote MIQCQPs in the following form:
@@
[…]
$x^{L}_{k}\leqslant\ x_{k}\leqslant\ x^{U}_{k}$ $\quad\forall\,k\in N,$
$x_{k}\in\ \mathbb{Z}$ $\quad\forall\,k\in I,$
where $N:=\{1,\dots,n\}$, $I\subseteq N$ is the index set of integer variables, $Q,A_{i}\in\mathbb{Q}^{n\times n}$ are symmetric matrices, $d,\ b_{i}\in\mathbb{Q}^{n}$, $c_{i}\in\mathbb{Q}$ for $i=1,\dots,m$, and $x^{L},x^{U}\in\bar{\mathbb{Q}}^{n}$ (with $\bar{\mathbb{Q}}:=\mathbb{Q}\cup\{\pm\infty\}$) are the lower and upper bounds of $x$. Note that both $Q$ and $A_{i}$ need not be positive semidefinite, hence nonconvex quadratic problems are allowed. If $I=\emptyset$, the problem reduces to a quadratically constrained quadratic program (QCQP). In the following, when discussing a single quadratic constraint, we may omit the constraint index $i$ for simplicity.
In this work, we present a framework for this class of problems with an emphasis on finding primal feasible solutions, which we developed in the context of the Mixed-Integer Programming workshop 2025  [^1] Details on the competition can be found at the following page:
@@
. The framework also provides dual bounds for a subset of MIQCQPs as a by-product of the primal search. Our approach is based on a branch-and-bound algorithm leveraging Frank-Wolfe to optimize continuous relaxations over the convex hull of mixed-integer feasible points, as proposed in the Boscia framework [21]. Importantly, our method can leverage mixed-integer linear optimization solvers to compute vertices of a relaxed feasible region and can naturally handle differentiable nonlinear terms in the objective function, a property we leverage to transform and relax constraints. Guided by the nonlinear relaxations, we fix continuous and integer variables and use other off-the-shelf solvers for the resulting small problems, adapting techniques from the large neighborhood search literature. A high-level overview of our framework is presented in Figure 1.
@@
The rest of the paper is organized as follows. Section 2 presents the core part of the framework consisting of the specialized branch-and-bound with relaxations solved with Frank-Wolfe. Section 3 details the transformations we operate on the original MIQCQP, modifying the set of constraints and the objective. Section 4 expands on the heuristics called during and after the branch-and-bound process. Section 5 presents computational results assessing the performance of the framework and the impact of its different components.
2 Boscia and Frank-Wolfe
@@
where $\mathcal{X}\subseteq\mathbb{R}^{n}$ is a convex, compact (polyhedral) set and $I\subseteq N$ is the set of integer variables.
@@
It operates a branch-and-bound scheme, utilizing Frank-Wolfe (FW) methods [11] implemented in the FrankWolfe.jl library [6, 7] to solve continuous subproblems. FW methods are projection-free first-order methods that only require oracle access to the objective function, its gradient and the feasible region. At iteration $t$, the classic FW algorithm, dubbed Vanilla Frank-Wolfe, calls the Linear Minimization Oracle (LMO) over the feasible region using the gradient of the current iterate $x_{t}$ as the objective function. The solution $v_{t}$ is used to generate the next descent direction $d_{t}=x_{t}-v_{t}$ and the iterate is updated as $x_{t+1}=x_{t}-\gamma_{t}d_{t}$ where $\gamma_{t}$ is the step-size. Thus, the iterate can be represented as a convex composition of the extreme points of the feasible region. The composition is referred to as the active set in FW context. Many FW variants explicitly use the active set to make progress. The default variant used in Boscia, and in our set-up, is the Blended Pairwise Conditional Gradient (BPCG) method [32]. For the step-size rule, we use the secant line-search step-size rule [20], which is particularly efficient in the context of quadratic problems.
@@
Boscia solves a Mixed-Integer Programming (MIP) problem as the LMO. By doing so, Boscia optimizes convex relaxations over the convex hull of mixed-integer feasible points instead of the continuous relaxation of the constraints, resulting in a much smaller branch-and-bound search and directly leveraging the MIP solver machinery, e.g., cutting planes, conflicts analysis, heuristics. This adjustment enables Boscia to inherently sample integer-feasible solutions, effectively embedding a heuristic search within the optimization process. Importantly, optimizing over the convex hull of integer-feasible solutions is performed by Boscia without an explicit algebraic description of this feasible region. In order to alleviate the cost of the multiple MIP solves, Boscia integrates lazified FW variants which reuse vertices of the feasible region computed by the MIP solver as much as possible, which can be performed without losing convergence guarantees; see, e.g., [12] for details on this lazification. For further lazification, Boscia holds on to vertices dropped by FW as they might become useful for nodes later down the branch. Hence, any integer feasible point is computed at most once. The active set is used during branching to facilitate warm-starting of the children nodes by splitting the active set of their parent.
@@
MIQCQPs do not directly fit into the convex mixed-integer optimization framework that Boscia is designed for, so several key modifications to the algorithm are needed. Since Boscia assumes linear constraints, the quadratic constraints must be reformulated to comply with this structure. We achieve this by employing a power penalty relaxation, ensuring that the solver can process these constraints while maintaining feasibility, see Section 3.2.
@@
Another key challenge is Boscia’s assumption of convexity, which does not always hold for the given MIQCQPs. To address this, we implement several modifications. First, we disable node pruning based on the lower bound in the branch-and-bound process, as the lower bound information may no longer be valid. Second, we instruct Boscia to ignore the standard FW lower bound, which is not applicable in our setting. Finally, we adjust Boscia’s solution storage mechanism. By default, it performs pre-sampling and retains only improving solutions, but this approach proved inadequate for quadratic constraints. Instead, we modify the solver to track all solutions encountered during the process, ensuring that valuable candidates are not prematurely discarded. Additionally, we added a callback mechanism that allows us to evaluate the solutions with respect to the original objective and lets us discard solutions which do not satisfy the quadratic constraints.
@@
These modifications allow Boscia to effectively handle the challenges posed by quadratic mixed-integer problems, making it a competitive approach for the Land-Doig MIP Computational Competition 2025. Even though the focus of the competition is on the primal side, we highlight that the developed solution framework is also very suited to derive high-quality relaxation bounds, since the constraint penalization and many transformations are exact reformulations or produce relaxations of the original problem.
@@ 3 Problem transformations and presolving
In this section, we present the problem transformations performed before executing the main algorithms. These presolving steps are crucial to obtain good performance on several classes of problems.
@@ 3.1 Propagation
Initially, we apply a propagation step to tighten the bounds of the variables. For linear constraints, we apply a simple activity-based bound strengthening technique and some specialized propagators for several classes of linear constraints [16]. For quadratic constraints, we experimented with McCormick linearizations [27], introducing auxiliary variables $z_{ij}$ to replace bilinear terms $x_{i}x_{j}$, and constructing McCormick envelopes. However, we found that incorporating these linearized constraints in the problem formulation, or even just propagating them, shows little to no performance improvement in practice.
@@ 3.2 Power penalty
[…]
for a real exponent $p>1$. A parameter $p=1$ would result in a nonsmooth nondifferentiable function, while $p=2$ would result in a term akin to that of augmented Lagrangian methods. We experimented with $p\in[1.2,1.8]$, trading off steepness of the gradient towards feasible points and smoothness of the continuous subproblems tackled by Frank-Wolfe.
@@ 3.3 Quadratic special structures
Complementarities
are constraints of the form $x_{i}x_{j}=0$. Although they are quadratic equalities, they are combinatorial in nature and can be treated as such with an additional binary variable $z$ controlling which of the two terms should be set to zero. In order to avoid relying on variable bounds for a big-M constraint, we reformulate the complementarity to a pair of indicator constraints, which we present here with the assumption that both variables are nonnegative:
$z=0\Rightarrow x_{i}\leqslant 0$
@@
Importantly, we do not reformulate complementarity quadratic constraints to Special Ordered Sets of type 1 (SOS1) constraints, since such constraint would not respect the assumptions from Boscia (see [22]).
Perspective constraints
are another type of special quadratic constraint from three variables: $x$ is a single variable, $w$ is the epigraph variable, and $z$ is a binary variable. The constraint is of the form:
@@
$\left\{(x,w,z)\in\mathrm{R}_{+}\times\mathrm{R}_{+}\times\{0,1\}\,:\,x^{2}\leqslant w,(1-z)x=0\right\},$
see, e.g., the seminal paper [17] on perspective functions in mixed-integer convex optimization and [19] and references therein on handling perspective functions in a nonlinear solver. The binary variable activates the continuous variable $x$, the epigraph variable $w$ is often a consequence of a nonlinear objective term converted to a constraint. When this is the case and the epigraph $w$ indeed only appears in the perspective constraint and in the objective, we transform the constraint back into a quadratic term $cx^{2}$ in the objective along with a big-M and/or indicator constraint explicitly tying $z$ and $x$. This transformation separates the nonsmooth perspective constraint that can lead to numerical challenges in a nonlinear setting (as reported in [19] among others) into a smooth quadratic term in the objective and the combinatorial “activation” structure in the constraints.

### v2

@@ Abstract
We propose a primal heuristic for quadratic mixed-integer problems. Our method extends the Boscia framework – originally a mixed-integer convex solver leveraging a Frank-Wolfe-based branch-and-bound approach – to address nonconvex quadratic objective functions and constraints. We reformulate nonlinear constraints, introduce preprocessing steps, and a suite of heuristics including rounding strategies, gradient-guided selection, and large neighborhood search techniques that exploit integer-feasible vertices generated during the Frank-Wolfe iterations. Computational results demonstrate the effectiveness of our method in solving challenging MIQCQPs, achieving improvements on QPLIB instances within minutes and winning first place in the Land-Doig MIP Computational Competition 2025.
@@ 1 Introduction
Mixed-Integer Quadratically Constrained Quadratic Problems (MIQCQPs) represent a broad category in optimization, capturing an array of applications in operations research and machine learning. These problems combine the combinatorial difficulties of mixed-integer structures with nonlinear nonconvex constraints. Despite this double difficulty, solution methods, algorithms, and solvers have been developed over the past decades for generic or specific forms of MIQCQPs, including Couenne [2], GloMIQO [32], ANTIGONE [33], BARON [38], and SCIP [8, 11, 9]. We refer the reader to [27, 28] for recent reviews of mixed-integer nonlinear optimization beyond quadratic functions. Recently, commercial mixed-integer linear solvers such as Xpress and Gurobi added capabilities to solve MIQCQPs to global optimality, showing the strong interest from application areas. We consider MIQCQPs in the following form:
@@
s.t. $x^{\top}A_{k}x+b_{k}^{\top}x+c_{k}\leqslant 0$ $\quad\forall\,k=1,\dots,m,$
$x^{L}_{i}\leqslant x_{i}\leqslant x^{U}_{i}$ $\quad\forall\,i\in N,$
$x_{i}\in\mathbb{Z}$ $\quad\forall\,i\in I,$
where $n,m\in\mathbb{N}$ are the number of variables and constraints, $N:=\{1,\dots,n\}$, $I\subseteq N$ is the index set of integer variables, $Q,A_{k}\in\mathbb{Q}^{n\times n}$ are symmetric matrices, $d,\ b_{k}\in\mathbb{Q}^{n}$, $c_{k}\in\mathbb{Q}$ for $k=1,\dots,m$, and $x^{L},x^{U}\in\bar{\mathbb{Q}}^{n}$ (with $\bar{\mathbb{Q}}:=\mathbb{Q}\cup\{\pm\infty\}$) are the lower and upper bounds of $x$. Note that both $Q$ and $A_{k}$ need not be positive semidefinite, hence nonconvex quadratic problems are allowed. If $I=\emptyset$, the problem reduces to a quadratically constrained quadratic program (QCQP). In the following, when discussing a single quadratic constraint, we may omit the constraint index $k$ for simplicity.
In this work, we present a framework for this class of problems with an emphasis on finding primal feasible solutions, which we developed in the context of the Land-Doig MIP Computational Competition 2025  [^1] Details on the competition can be found at the following page:
@@
. The framework also provides dual bounds for a subset of MIQCQPs as a by-product of the primal search. Our approach is based on a branch-and-bound algorithm leveraging Frank-Wolfe to optimize continuous relaxations over the convex hull of mixed-integer feasible points, as proposed in the Boscia framework [24]. Importantly, our method uses mixed-integer linear optimization solvers to compute vertices of a relaxed feasible region and can naturally handle differentiable nonlinear terms in the objective function, a property we leverage to transform and relax constraints. Guided by the nonlinear relaxations, we fix continuous and integer variables and use other off-the-shelf solvers for the resulting smaller problems, adapting techniques from the large neighborhood search literature. Our framework can thus be contrasted with most MINLP and MIQCQP solvers mentioned above which primarily construct polyhedral relaxations, with a few exceptions such as BARON. A high-level overview of our framework is presented in Figure 1.
@@
The rest of the paper is organized as follows. Section 2 presents the core part of the framework consisting of the specialized branch-and-bound with relaxations solved with Frank-Wolfe. Section 3 details the transformations we operate on the original MIQCQP, modifying the set of constraints and the objective function. Section 4 expands on the heuristics called during and after the branch-and-bound process. Section 5 presents computational results assessing the performance of the framework and the impact of its different components.
2 Boscia: Integer Frank-Wolfe
@@
where $\mathcal{X}\subseteq\mathbb{R}^{n}$ is a convex, compact (polyhedral) set, $I\subseteq N$ is the set of integer variables and $f$ is a convex, continuously differentiable function and its gradient is Lipschitz continuous with constant $L$.
@@
It operates a branch-and-bound scheme, utilizing Frank-Wolfe (FW) methods [13] implemented in the FrankWolfe.jl library [6, 7] to solve continuous subproblems. FW methods are projection-free first-order methods that only require oracle access to the objective function, its gradient, and the feasible region. At iteration $t$, the classic FW algorithm, dubbed Vanilla Frank-Wolfe, calls the Linear Minimization Oracle (LMO) over the feasible region using the gradient of the current iterate $x^{t}$ as the objective function. The solution $v^{t}$ is used to generate the next descent direction $d^{t}=x^{t}-v^{t}$, and the iterate is updated as $x^{t+1}=x^{t}+\gamma^{t}d^{t}$ where $\gamma^{t}$ is the step size. Thus, the iterate can be represented as a convex combination of the extreme points of the feasible region. This combination is referred to as the active set in the FW context. Many FW variants explicitly use the active set to make progress. The Corrective Frank-Wolfe (CFW) algorithm [22] shown in Algorithms 1 and 2 presents a framework for all active set based variants of FW.
In lines 46, the so-called away vertex $a^{t}$, the vertex in the active set with the largest inner product with the gradient, the local FW $s^{t}$, vertex of the active set with the smallest inner product with the gradient, and the global FW vertex $v^{t}$ are computed. In line 7, $a^{t}$ and $s^{t}$ are used to compute the local gap. If it is sufficiently large, the algorithm proceeds to optimize the active set using the Corrective Step algorithm 2. Note that the constant $L$ is the Lipschitz constant of the gradient of the objective function and $D$ is the diameter of the feasible region $\mathcal{X}$. The specific corrective step depends on the variant of FW used. The default variant used in Boscia, and in our set-up, is the Blended Pairwise Conditional Gradient (BPCG) method [36]. The corrective step in this case shifts weight from the away vertex $a^{t}$ to the local FW vertex $s^{t}$, computed via linesearch. If all of the weight is shifted to the local FW vertex, the away vertex is dropped (drop step), otherwise the descent step is performed with descent direction $d=a^{t}-s^{t}$. As step-size rule, we use the secant line-search step-size rule [23], which is particularly efficient in the context of quadratic problems.
Algorithm 1 Corrective Frank-Wolfe (CFW) [22]
1:convex, smooth function $f$, start point $x^{0}\in V(\mathcal{X})$ (vertex of $\mathcal{X}$).
[…]
3:for $t=0$ to T1T-1 do
4:  $a^{t}\leftarrow\arg\max_{v\in S^{t}}\langle\nabla f(x^{t}),v\rangle$ $\triangleright$ away vertex
5:  $s^{t}\leftarrow\arg\min_{v\in S^{t}}\langle\nabla f(x^{t}),v\rangle$ $\triangleright$ local FW
6:  $v^{t}\leftarrow\arg\min_{v\in V(\mathcal{X})}\langle\nabla f(x^{t}),v\rangle$ $\triangleright$ global FW
7:  if $\langle\nabla f(x^{t}),a^{t}-s^{t}\rangle\geqslant\langle\nabla f(x^{t}),x^{t}-v^{t}\rangle$ then
8:   $x^{t+1},S^{t+1}\leftarrow\textsc{CorrectiveStep}(S^{t},x^{t},a^{t},s^{t})$
9:  else
10:   $\gamma^{t}\leftarrow\arg\min_{\gamma\in[0,1]}f(x^{t}-\gamma(x^{t}-v^{t}))$
11:   $x^{t+1}\leftarrow x^{t}-\gamma^{t}(x^{t}-v^{t})$
12:   $S^{t+1}\leftarrow S^{t}\cup\{v^{t}\}$
13:  end if
14:end for
Algorithm 2 Corrective Step($S,x,a,s$) [22]
[…]
@@
Boscia solves a Mixed-Integer Programming (MIP) problem as the LMO. By doing so, Boscia optimizes convex relaxations over the convex hull of mixed-integer feasible points instead of the continuous relaxation of the constraints, resulting in a much smaller branch-and-bound search tree and directly leveraging the MIP solver machinery, e.g., cutting planes, conflict analysis, and heuristics. This adjustment enables Boscia to inherently sample integer-feasible solutions, effectively embedding a heuristic search within the optimization process. Importantly, optimizing over the convex hull of integer-feasible solutions is performed by Boscia without an explicit algebraic description of this feasible region. In order to alleviate the cost of the multiple MIP solves, Boscia integrates lazified FW variants which reuse vertices of the feasible region computed by the MIP solver as much as possible, which can be performed without losing convergence guarantees; see, e.g., [14] for details on this lazification. For further lazification, Boscia holds on to vertices dropped by FW as they might become useful for nodes later down the branch. Hence, any integer-feasible point is computed at most once. The active set is used during branching to facilitate warm-starting of the children nodes by splitting the active set of their parent. Since all vertices are integer-feasible, any vertex $v$ of the active set must satisfy either $v_{i}\leqslant\lfloor\bar{x}_{i}\rfloor$ or $v_{i}\geqslant\lceil\bar{x}_{i}\rceil$ where $\bar{x}$ denotes the fractional node solution and $i$ the branching index. This property is used to split the active set into two parts during branching.
@@
MIQCQPs do not directly fit into the convex mixed-integer optimization framework that Boscia is designed for, therefore several key modifications to the algorithm are needed. Since Boscia assumes linear constraints, the quadratic constraints must be reformulated to comply with this structure. We achieve this by employing a power penalty relaxation, ensuring that the solver can process these constraints while maintaining feasibility (see Section 3.2).
@@
Another key challenge is Boscia’s assumption of convexity, which does not always hold for the given MIQCQPs. To address this, we implement several modifications. First, we disable node pruning based on the lower bound in the branch-and-bound process, as the lower bound information may no longer be valid. Second, we instruct Boscia to ignore the standard FW lower bound, which is not applicable in our setting. Finally, we adjust Boscia’s solution storage mechanism. By default, it performs pre-sampling and retains only improving solutions, but this approach proved inadequate for quadratic constraints. Instead, we modify the solver to track all solutions encountered during the process, ensuring that valuable candidates are not prematurely discarded. Additionally, we added a callback mechanism that allows us to evaluate the solutions with respect to the original objective and lets us discard solutions which do not satisfy the quadratic constraints. When Boscia finds a new solution, we check if it satisfies the quadratic constraints and discard it if it does not. Furthermore, the solution is only added to the list of solutions if it improves upon the current best solution.
Because we are primarily interested in finding solutions, we do not let Boscia solve to completion. Instead, we provide a node and time limit to the solver. If the node limit is reached and there is still time remaining, we restart Boscia with a new starting point. We compute the new start point via the LMO by using either the gradient of the current iterate or a random direction. In this way, we aim to explore as much of the feasible region as possible.
To test multiple problem modifications, we employ a parallelization strategy as it proved to be computationally inefficient to change problem parameters between restarts. For the convexification schemes, for example, we tried shifting different percentages of the spectrum of the Hessian into the positive orthant, see Section 3.4 for details. For problems with quadratic constraints, we tried different power penalty parameters, see Section 3.2 for details. Note that the restarts are performed in each thread.
@@
These modifications allow Boscia to effectively handle the challenges posed by quadratic mixed-integer problems, making it a competitive approach for the Land-Doig MIP Computational Competition 2025. Even though the focus of the competition is on the primal side, we highlight that the developed solution framework is also suited to derive high-quality relaxation bounds, since the constraint penalization and many transformations are exact reformulations or produce relaxations of the original problem.
@@ 3 Problem transformations and presolving
In this section, we present the problem transformations performed before executing the main algorithms. These presolving steps are crucial to obtain good performance on several problem classes.
@@ 3.1 Propagation
Initially, we apply a propagation step to tighten the bounds of the variables. For linear constraints, we apply a simple activity-based bound strengthening technique and some specialized propagators for several classes of linear constraints [18]. For quadratic constraints, we experimented with McCormick linearizations [30], introducing auxiliary variables $z_{ij}$ to replace bilinear terms $x_{i}x_{j}$, and constructing McCormick envelopes. However, we found that incorporating these linearized constraints in the problem formulation, or even just propagating them, showed little to no performance improvement in practice.
@@ 3.2 Power penalty
$f(x)+\mu\cdot{\max\{x^{\top}Ax+b^{\top}x+c,0\}}^{p}$
for a penalty parameter $\mu>0$ and a real exponent $p>1$. A parameter $p=1$ would result in a nonsmooth, nondifferentiable function, while $p=2$ would result in a term akin to that of augmented Lagrangian methods. We experimented with $p\in[1.2,1.8]$, trading off the steepness of the gradient towards feasible points and the smoothness of the continuous subproblems tackled by Frank-Wolfe.
@@ 3.3 Quadratic special structures
Complementarities.
They are constraints of the form $x_{i}x_{j}=0$. Although they are quadratic equalities, they are combinatorial in nature and can be treated as such with an additional binary variable $z$ controlling which of the two terms should be set to zero. In order to avoid relying on variable bounds for a big-M constraint, we reformulate the complementarity to a pair of indicator constraints, which we present here with the assumption that both variables are nonnegative:
$z=0\Rightarrow x_{i}\leqslant 0,$
@@
Importantly, we do not reformulate complementarity quadratic constraints to Special Ordered Sets of type 1 (SOS1) constraints, since such constraints would not respect the assumptions from Boscia (see [25]).
Perspective constraints.
They are another type of special quadratic constraint from three variables: $x$ is a single variable, $w$ is the epigraph variable, and $z$ is a binary variable. The constraint is of the form:
@@
$\left\{(x,w,z)\in\mathbb{R}_{+}\times\mathbb{R}_{+}\times\{0,1\}\,:\,x^{2}\leqslant w,(1-z)x=0\right\},$
see, e.g., the seminal paper [19] on perspective functions in mixed-integer convex optimization and [21] and references therein on handling perspective functions in a nonlinear solver. The binary variable activates the continuous variable $x$; the epigraph variable $w$ is often a consequence of a nonlinear objective term converted to a constraint. When this is the case and the epigraph $w$ indeed only appears in the perspective constraint and in the objective, we transform the constraint back into a quadratic term $cx^{2}$ in the objective along with a big-M and/or indicator constraint explicitly tying $z$ and $x$. This transformation separates the nonsmooth perspective constraint that can lead to numerical challenges in a nonlinear setting (as reported in [21] among others) into a smooth quadratic term in the objective and the combinatorial “activation” structure in the constraints.